# Detecting Masquerade Attacks in Controller Area Networks Using Graph Machine Learning

### v1

@@ VI Discussion
This study introduces a novel approach for detecting masquerade attacks in CAN using graph ML. By combining graph embeddings with time series features, our method demonstrates significant improvements in detection capabilities across various masquerade attack scenarios. Our results consistently show that this combined approach outperforms the graph embeddings-only method across all attack types examined in the ROAD dataset. The improvement in performance is statistically significant, as evidenced by the Mann-Whitney $\mathrm{U}$ and $\mathrm{KS}$ tests (see Tables III and IV). Additionally, our exploration of sample-based windows reveals promising results, particularly with smaller $\omega$ values, suggesting that fine-grained analysis of CAN messages can even capture subtle anomalies.
The enhanced performance of our combined approach emphasizes the importance of incorporating both structural and temporal information in CAN IDS. By leveraging graph embeddings, we capture the complex interactions between different ECUs, while the addition of time series features allows us to model the temporal dynamics of CAN signals. This comprehensive view enables our system to detect subtle deviations that are characteristic of sophisticated masquerade attacks. Moreover, our experimental results indicate that the integration of graph embeddings with time series features optimizes the detection parameters, reducing the required $\omega$ for optimal detection by 38.8% while increasing the optimal $\delta$ by 50%. This adjustment supports more efficient anomaly detection, crucial for timely responses in vehicular networks. The effectiveness of smaller $\omega$ values coupled with larger $\delta$ values suggests that detailed analyses of CAN messages can reveal anomalous behavior patterns while being performed less frequently, striking a balance between granularity and efficiency. This approach could be particularly beneficial in resource-constrained vehicular environments, with important implications for the design of potential real-time IDS.
@@
Lack of Validation in a Real Environment: Our study is simulation-based and lacks validation in a real vehicle environment. Real-world factors such as network latency, sensor noise, and varying driving conditions could potentially impact the performance of our system. We partitioned logs of already collected CAN captures to simulate a streaming environment of CAN data using windows. Future work should involve empirical validation on a moving vehicle, including the necessary hardware and performance tuning involved to do so.
Evaluation Based Solely on ROAD Dataset: Our evaluation is based solely on the ROAD dataset [15], which, although comprehensive, may not capture the full range of potential attack scenarios in real-world settings. This limitation suggests that future studies should incorporate a variety of datasets, including those with more recent and diverse attack scenarios, to enhance the robustness and generalizability of the findings.
Computational Complexity: The computational complexity of our approach, particularly in the graph embedding phase, may present challenges for real-time implementation in resource-constrained vehicular environments. Further optimization may be necessary for practical deployment. Future research should focus on developing more efficient algorithms and exploring hardware acceleration to make the approach viable for real-time applications.
Use of Default Parameter Values in Detection Algorithms: We used the same default parameter values as the original algorithms, such as the number of trees and maximum depth for the Random Forest classifier, and similar default settings for the XGBoost algorithm. While this ensures consistency in evaluation, it may not represent the optimal configuration for each algorithm in the context of our specific application. This limitation underscores the importance of performing extensive hyperparameter tuning in future work to identify the best settings for each algorithm to maximize their detection performance.
@@ VII Conclusion
This paper introduces a novel framework for detecting masquerade attacks in CANs using graph machine learning. Our approach uniquely combines graph embeddings with time series analysis, significantly enhancing detection capabilities across various attack patterns. By representing CAN frames as MSGs and enriching nodes with statistical attributes from time series data, we achieve a more comprehensive view of network behavior. This integration allows for the detection of subtle anomalies characteristic of sophisticated masquerade attacks, outperforming methods based solely on graph embeddings across all tested attack types in the ROAD dataset [15]. Our method demonstrates statistically significant improvements (see Tables III and IV), particularly in detecting challenging attacks like the max speedometer attack, which has proven elusive in previous studies [60].
Looking forward, we plan to continue expanding this research in optimizing our approach for real-time implementation in resource-constrained vehicular environments. This could involve exploring more efficient graph neural network architectures, such as those proposed by Zhang et al. [23]. Additionally, extending the evaluation to a broader range of datasets and attack scenarios will be crucial for validating the generalizability of our method. As vehicles become increasingly connected and autonomous, our work contributes to the critical field of automotive cybersecurity, paving the way for more robust and adaptive IDS in future vehicles and cyber-physical systems.

### v2

@@ VI Discussion
This paper presents a novel approach for detecting masquerade attacks in CAN using graph ML. By combining graph embeddings with time series features, our method significantly improves detection performance across all masquerade attack types in the ROAD dataset. These gains are statistically validated using the Mann-Whitney $\mathrm{U}$ and $\mathrm{KS}$ tests (see Tables VII and IX). For instance, detection of the max_engine_coolant_attack—previously difficult to detect due to weak signal deviation [moriano2024benchmarkingunsupervisedonlineids]—was notably improved in the enhanced setting, demonstrating stronger anomaly sensitivity. Our evaluation also shows that using sample-based windows—particularly with smaller $\omega$ values—enables the detection of subtle anomalies via fine-grained analysis.
What specific deviations does the model capture?
Our hybrid framework captures subtle, cumulative deviations that even sophisticated spoofing cannot perfectly hide. We leverage two complementary domains:
  • Structural deviations (via MSGs and graph embeddings): Masquerade attacks involve an attacker suppressing a legitimate ECU and impersonating it. Attackers cannot perfectly replicate intricate real-time inter-message dependencies, message orderings, and complex sequential patterns (graph motifs) characteristic of benign ECU communication. Prior work demonstrates motif frequency distributions as sensitive indicators of anomalies [milo2002network]. Graph-based approaches effectively capture these structural anomalies in CAN traffic. Our approach leverages node2vec to embed these deviations, extracting low-dimensional representations that preserve the complex topology of MSGs.
  • Temporal and content deviations (via timeseries features): Masquerade attacks can introduce subtle inconsistencies in physical signal values or precise timing patterns, even if overall frequency is matched. Decoded signal values may show anomalies in statistical properties (mean, standard deviation) or inter-signal correlations, which are stable in normal driving but deviate under spoofing [zhang2020cansec]. Our node annotation technique enriches CAN ID nodes with these statistics, enabling detection of nuanced temporal and content-level inconsistencies. Similar approaches prove effective for high-frequency signal perturbations in automotive IDS [Zhang2023, moriano2024benchmarkingunsupervisedonlineids].
The strength of our hybrid approach lies in its ability to fuse these structural graph embeddings with the temporal signal statistics. A deviation missed by one modality might be detected by the other, providing a robust, multi-faceted detection capability against attacks that blend seamlessly into normal traffic.
The effectiveness of our approach highlights the value of combining structural and temporal insights in CAN IDS. Graph embeddings capture ECU interactions, while time series features reflect evolving signal patterns. This dual analysis improves detection and allows more efficient sliding window configuration—reducing the optimal offset $\delta_{t}$ by 38.8% (9.8 s to 6 s) and increasing the window length $\omega_{t}$ by 50% (3.2 s to 4.8 s). These adjustments enhance timely detection and emphasize the importance of securing CAN-based systems as vehicle architectures evolve.
While automotive Ethernet adoption grows (projected to reach 11.6 billion US dollars by 2032 [GMI2024], it primarily serves high-bandwidth domains like ADAS and infotainment (38% market share in 2023). However, CAN remains critical for safety subsystems (e.g., engine, braking) due to its role in hybrid architectures and persistent vulnerabilities. Over 90% of modern vehicles use hybrid CAN/Ethernet networks, with CAN handling safety-critical low-speed data (powertrain/chassis) while Ethernet supplements infotainment [GMI2024]. CAN’s simplicity and cost ensure its ubiquity in existing fleets, with Ethernet adoption concentrated in high-end models. Despite Ethernet’s growth, CAN lacks encryption/authentication, making it vulnerable to attacks like brake disablement via OBD-II injection [SealingTech2024]. These risks persist even in hybrid architectures, as compromised Ethernet gateways can inject malicious CAN messages. Our work addresses these gaps with lightweight CAN-specific detection mechanisms. CAN security remains foundational, as even Ethernet-reliant systems interface with CAN for critical functions.
Our work distinguishes itself from prior methods by integrating graph-based and time series analysis to address the subtle nature of masquerade attacks. While existing approaches focus on single-modality detection, our hybrid framework captures both structural and temporal anomalies, enabling the detection of sophisticated intrusions that traditional IDS might miss. This dual approach not only improves detection accuracy but also provides a scalable solution adaptable to the dynamic nature of modern vehicular networks.
Our parameter choices are empirically grounded in the ROAD dataset’s attack profiles, balancing accuracy and practicality. While shorter windows may suit other threat models, masquerade detection necessitates analyzing sustained patterns. For safety-critical functions requiring sub-second detection (e.g., brake disengagement), adaptive parameter tuning and hardware-specific optimizations will be explored in future work. While our approach demonstrates promising detection capabilities, its real-time deployment in resource-constrained vehicular environments requires further validation on embedded platforms (e.g., ARM-based boards). Our current results, based on realistic CAN logs generated on a physical vehicle, focus on validating the detection methodology rather than finalizing an on-vehicle implementation.
While cryptographic measures such as encryption and MACs are effective defenses against masquerade attacks, our anomaly detection framework provides an additional layer of security. By detecting subtle deviations in CAN traffic patterns, our approach complements cryptographic defenses, ensuring robust protection against sophisticated attacks, including those that bypass cryptographic approaches.
@@
Lack of Validation in a Real Environment: Our study is simulation-based and lacks validation in real vehicle environments, where factors like latency, sensor noise, and varying driving conditions could impact performance. Future work should involve empirical validation on a moving vehicle, including hardware integration and performance tuning.
Evaluation Based Solely on ROAD Dataset: While comprehensive, the ROAD dataset [verma2022addressing] may not capture the full range of real-world attack scenarios. Future studies should incorporate additional datasets with diverse and recent attack patterns to improve robustness and generalizability.
Computational Complexity: The graph embedding phase may introduce computational overhead. Future work should explore algorithmic optimization and hardware acceleration for improved efficiency.
Use of Default Parameter Values in Detection Algorithms: This study relied on default parameters in detection algorithms, which may not yield optimal performance. Future research should involve extensive hyperparameter tuning to maximize detection accuracy in application-specific contexts.
Limits of Detectability: No system is impervious to highly sophisticated attackers. The detectability limits for our framework are as follows:
  • Perfect mimicry: If an attacker can replicate the full message transition topology and multivariate signal relationships of a legitimate ECU—including physical-layer jitter and temporal correlations—the detection margin narrows. However, this level of mimicry requires access to proprietary ECU timing, bus arbitration, and signal generation logic, which is extremely difficult without insider knowledge [Miller2014RemoteAttack].
  • Short or bursty attacks: Transient attacks that operate below the minimum window size or inject very few anomalous transitions may evade detection. This is a known limitation in time-window-based IDS designs [Jedh2021].
  • Overlapping statistical distributions: In rare cases, spoofed messages may fall within the natural variability of benign traffic, making them indistinguishable even in a fused feature space. Defense against this class of attack may require incorporating physical-layer fingerprints [cho2016fingerprinting] or entropy-based signal modeling [Marchetti2017].
@@ VII Conclusion
This paper presents a novel framework for detecting masquerade attacks in CANs by integrating graph ML with time series analysis. By representing CAN frames as MSGs and enriching each node with contextual statistical features, the proposed approach captures both the structural and temporal dynamics of in-vehicle communication. Experiments on the ROAD dataset [verma2022addressing] show that this hybrid model consistently outperforms graph-only baselines across all tested masquerade attack types. These improvements are statistically validated through the Mann-Whitney $\mathrm{U}$ and Kolmogorov-Smirnov tests (see Tables VII and IX), confirming the robustness of our detection capabilities across various attack scenarios.
Future work will explore deploying this framework in online detection settings, investigate lightweight graph-based architectures such as those introduced by Zhang et al. [Zhang2023], and assess its adaptability across diverse datasets. We also aim to explore integration with foundation models for enhancing generalization in evolving vehicular environments.
@@ Acknowledgments
[Uncaptioned image] William Marfo (Member, IEEE) received the Ph.D. degree in Computer Science from the University of Texas at El Paso, USA. He is currently an Associate Specialist and Cyberinfrastructure Software Engineer with the Cascadia Region Earthquake Science Center (CRESCENT) at the University of California, Davis, USA. His research interests include machine learning, high-performance computing, and cybersecurity for complex computational systems. He develops scalable software and AI/ML models for large-scale seismic data processing and interdisciplinary earthquake science. He previously conducted research at U.S. Department of Energy national laboratories and was named a 2025 Trusted CI Fellow by the US National Science Foundation (NSF) Cybersecurity Center of Excellence. Dr. Marfo is a member of ACM, SIAM, AMS, and ISACA.
[Uncaptioned image] Pablo Moriano (Senior Member, IEEE) received the B.S. and M.S. degrees in electrical engineering from Pontificia Universidad Javeriana, Colombia, and the M.S. and Ph.D. degrees in informatics from Indiana University Bloomington, Bloomington, IN, USA. He is a Staff Scientist with the Computer Science and Mathematics Division, Oak Ridge National Laboratory, Oak Ridge, TN, USA. His research interests include data science, network science, and cybersecurity, with a focus on the design of AI and statistical methods for anomaly detection in large-scale networked and cyber-physical systems. Applications of his work include intrusion detection, insider-threat analytics, and Internet-routing security. Dr. Moriano is also a Senior Member of the ACM and a member of SIAM.
[Uncaptioned image] Deepak Tosh (Senior Member, IEEE) is an associate professor of Computer Science at the University of Texas at El Paso. His research focuses on addressing various multi-disciplinary networking and cybersecurity challenges associated with critical national infrastructures, network virtualization, Industrial Internet of Things (IIoT), space cyber-physical systems, Blockchain, and tactical networking. His research has been supported by Department of Energy (DOE) and National Science Foundation (NSF) grants. He has authored/co-authored more than 90 peer-reviewed conference papers, book chapters, and journal papers. Two of his research works on Blockchain and data provenance received the “Top 50 Blockchain Papers in 2018” award at BlockchainConnect Conference, 2019. He is also a recipient of the prestigious NSF CAREER award, 2023.
[Uncaptioned image] Shirley V. Moore (Member, IEEE) received a B.A. degree in mathematics and chemistry from Indiana University, an M.Ed. degree from the University of Illinois, an M.S. in mathematics from Wichita State University, and M.S. and Ph.D. degrees in Computer Sciences from Purdue University. She is an Associate Professor in the Computer Science Department at the University of Texas at El Paso, USA. Her research is in parallel and high-performance computing, including distributed deep learning. Her focus is on performance optimization methodologies and tools. She is a Senior Member of ACM.